# Character Theory for Semilinear Representations

### v2

@@ Abstract.
Let $G$ be a group acting on a field $L$, and suppose that $L/L^{G}$ is a finite extension. We show that the irreducible semilinear representations of $G$ over $L$ can be completely described in terms of irreducible linear representations of $H$, the kernel of the map $G\rightarrow\operatorname{Aut}(L)$. When $G$ is finite and $|G|\in L^{\times}$ this provides a character theory for semilinear representations of $G$ over $L$, which recovers ordinary character theory when the action of $G$ on $L$ is trivial.
@@ 1. Introduction
$\rho(g)(\lambda\cdot v)=\sigma_{g}(\lambda)\cdot\rho(v)\ \mbox{ for all }\ g\in G,\lambda\in L,v\in V.$
@@ 1.2. The Extension Problem
Similarly, if $W$ is the non-trivial $1$-dimensional representation of $C_{2}$ over ${\mathbb{Q}}(\sqrt{d})$, then to give a semilinear representation of $C_{4}$ over ${\mathbb{Q}}(\sqrt{d})$ extending $W$ is the same as giving an $a\in{\mathbb{Q}}(\sqrt{d})$ with
@@ 1.5. Group Action
1.6. Irreducible Representations
We now restrict our attention to $\operatorname{Irr}_{L}^{\rtimes}(G)$, the set of isomorphism classes of irreducible semilinear representations of $G$ over $L$. We similarly write $\operatorname{Irr}_{L}(H)$ for the set of isomorphism classes of irreducible representations of $H$ over $L$.
Our next result describes $\operatorname{Irr}_{L}^{\rtimes}(G)$ in terms of $\operatorname{Irr}_{L}(H)$ and its action of $\operatorname{Gal}(L/K)$. To ease notation, we set $\Gamma\coloneqq\operatorname{Gal}(L/K)$, and for $W\in\operatorname{Irr}_{L}(H)$ write $\Gamma_{W}$ for the stabiliser of $W$ in $\Gamma$. Along with the restriction functor
@@ 1.6. Irreducible Representations
Suppose that $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$ and $W\in\operatorname{Irr}_{L}(H)$, where $W$ is an irreducible submodule of $V|_{H}$, or equivalently $V$ is an irreducible submodule of $\operatorname{Ind}_{H}^{G}(W)$. Then:
@@
    for some integer $m(V)\geq 1$.
  • In particular, $V|_{H}$ and $\operatorname{Ind}_{H}^{G}(W)$ are semisimple, and $m(V)\mid|\Gamma_{W}|$.
  • $\operatorname{Ind}_{H}^{G}(W)\cong\operatorname{Ind}_{H}^{G}(\gamma*W)$ for any $\gamma\in\operatorname{Gal}(L/K)$.
  • Any $W\in\operatorname{Irr}_{L}(H)$ is contained in $V|_{H}$ for a unique $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$.
  • This defines a bijective correspondence
    $\operatorname{Irr}_{L}^{\rtimes}(G)\leftrightarrow\operatorname{Irr}_{L}(H)/\hskip 1.00006pt\Gamma.$
In particular, this gives a complete description of $\operatorname{Irr}_{L}^{\rtimes}(G)$, up to a knowledge of the numbers $m(V)$: for any orbit of the action of $\operatorname{Gal}(L/K)$ on $\operatorname{Irr}_{L}(H)$, $m(V)$ copies of the orbit sum extends uniquely to an irreducible semilinear representation of $G$ over $L$, and all elements of $\operatorname{Irr}_{L}^{\rtimes}(G)$ arise this way.
@@ 1.7. Semilinear Schur Index
For $W\in\operatorname{Irr}_{L}(H)$, set $m_{K}^{L}(W)\coloneqq m(V)$ for the unique corresponding $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$.
We refer to $m_{K}^{L}(W)$ as the semilinear Schur index of $W$, as these generalise the classical Schur indices in an appropriate sense (see Section 10). Rephrasing Theorem B, these measure the failure for a $\Gamma$-fixed $W\in\operatorname{Irr}_{L}(H)$ to extend to a semilinear representation.
@@
$W\in\operatorname{Irr}_{L}(H)$ extends to a semilinear representation of $G$ over $L$ if and only if $\Gamma$ fixes $W$ and $m_{K}^{L}(W)=1$.
In Section 9 we establish the basic properties of the $m_{K}^{L}(W)$, including the fact that all $m_{K}^{L}(W)=1$ whenever $K$ is a finite field. In general, much like their classical counterparts, the numbers $m_{K}^{L}(W)$ are hard to determine, and often contain arithmetic information. For example, for $W$ as in Example 1.1, $m_{K}^{L}(W)$ is either $1$ or $2$, and $m_{K}^{L}(W)$ is $1$ if and only if $x^{2}-dy^{2}=-1$ has a solution over $\mathbb{Q}$}.
@@ 1.8. Character Theory
When $G$ is finite and $|G|$ is invertible in $L$, character theory provides a powerful and computationally efficient tool for completely describing the category $\operatorname{\mathbf{Rep}}_{L}(G)$.
@@
Suppose that $G$ is finite and $|G|$ is invertible in $L$. If $V,W\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, then:
@@
Another important consequence of the assumption that $G$ is finite and $|G|$ is invertible in $L$ is that $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ is semisimple with finitely many simple objects. This means that the description of $\operatorname{Irr}_{L}^{\rtimes}(G)$ provided by Theorem B suffices to describe the entire category $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, up to a knowledge of the semilinear Schur indices $m_{K}^{L}(W)$. For example, this gives a complete description of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ when $K$ is a finite field, as in this case all $m_{K}^{L}(W)=1$.
As an example of how this works, we give a complete classification of the semilinear representations of $S_{3}$ acting on a general degree two Galois extension $L/K$ via the natural projection $S_{3}\twoheadrightarrow\operatorname{Gal}(L/K)$. This is first done by hand in Section 5, with explicit semilinear matrix computations. We then return to this example later in Section 13 to recover this classification more conceptually using the theory provided by Theorems A and B.
@@ 1.10. Determination of Semilinear Schur Indices
In light of Theorem B, to understand $\operatorname{Irr}_{L}^{\rtimes}(G)$, one is reduced to understanding the semilinear Schur indices of $W\in\operatorname{Irr}_{L}(H)$. This is hard in general.
@@ 1.10.1. Other Associated Linear Categories
One natural way one might try to glean information about the semilinear Schur indices $m_{K}^{L}(W)$ is to use other linear representation categories associated with $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$. Theorem B and the theory developed above can be viewed as a partial description of $\operatorname{Irr}_{L}^{\rtimes}(G)$ in terms of $\operatorname{Irr}_{L}(H)$, but there is another linear representation category that one might naturally consider, namely $\operatorname{\mathbf{Rep}}_{K}(G)$.
@@ 1.11. Structure of the Paper
In Section 4 we study those semilinear representations arising from matrix representations defined over $K$, and give an extended example of how these can be used to describe all semilinear representations of $G$ over $L$ in Section 5, which also serves to demonstrate how one works with semilinear representations explicitly.
@@
In Section 7 we show that the semilinear representation category $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, for $L$ a potentially disconnected finite Galois extension of $K$, is equivalent to $\operatorname{\mathbf{Rep}}_{F}^{\rtimes}(G_{e})$, where $F$ is the field defined by a connected component $\operatorname{Spec}(L)$, and $G_{e}$ is the stabiliser. Aside from this reducing general semilinear representation categories to the case when $L$ is a field, this result is crucially used in Section 8 to prove Theorems A and B.
In Section 8, we prove Theorems A and B. The key idea is to interpret the functors $(-)|_{H}$ and $\operatorname{Ind}_{H}^{G}$ between $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ and $\operatorname{\mathbf{Rep}}_{L}(H)$ as certain honest base change functors for the field extension $L/K$. This allows us to port many properties of base change to our setting, such as the preservation of semisimplicity, and how the functors $(-)|_{H}$ and $\operatorname{Ind}_{H}^{G}$ compose. Theorem A follows from Corollary 8.6, Remark 2.13, Remark 3.5, and Remark 8.9, and Theorem B follows from Remark 3.5 and Theorem 8.20. We also relate the integers $m(V)$ to the $K$-division algebras $\operatorname{End}_{L\rtimes G}(V)$ in Proposition 8.28.
In Section 9 we define the semilinear Schur index of an object of $\operatorname{Irr}_{L}(H)$, and establish its basic properties, and in Section 10 we explain how the semilinear Schur indices specialise to the (classical) Schur indices in the case when $H\hookrightarrow G$ is split. We also highlight which properties the classical Schur indices satisfy which are not satisfied by all semilinear Schur indices.
In Section 11 we study the splitting behaviour of the $K$-algebras $\operatorname{End}_{L\rtimes G}(V)$ for $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$, and consider base change for semilinear representations. When $K$ is a number field we use the local-global principle for $\operatorname{Br}(K)$ to prove a local-global principle for semilinear representations.
@@ 2. Semilinear Representations
$\rho(g)(\lambda\cdot v)=\sigma_{g}(\lambda)\cdot\rho(v)\ \mbox{ for all }\ g\in G,\lambda\in R,v\in V.$
@@
Definition 2.3.
$\rho(g)(w)\in W\ \mbox{ for all }\ g\in G,w\in W.$
We say that $V$ is irreducible if $V$ is non-zero and 0 and $V$ are the only subrepresentations of $V$.
We write $\operatorname{Irr}_{R}^{\hskip 0.81949pt\sigma}(G)$ for the set of isomorphism classes of irreducible objects of $\operatorname{\mathbf{Rep}}_{R}^{\hskip 0.81949pt\sigma}(G)$.
@@ 2.1.2. Semilinear Matrix Representations
Let $n\coloneqq\dim_{L}(V)$, and fix an basis of the $R$-module $V$. For any $g\in G$, write $A_{g}\in M_{n\times n}(R)$ for the matrix of $\rho(g)\colon{}^{\sigma_{g}}V\rightarrow V$. The compatibility relations above in matrix form become
@@
We can give matrix descriptions of the constructions of Examples 2.4 and 2.5. The trivial representation has matrix representation by $(I_{1})_{g\in G}$, for $I_{1}$ the identity of $M_{1}(R)$, and if $V$ and $W$ in $\operatorname{\mathbf{Rep}}_{R}^{\hskip 0.81949pt\sigma}(G)$ have matrix representations $(A_{g})_{g\in G}$ and $(B_{g})_{g\in G}$ respectively, then
@@ 2.2. Semilinear Representations and Twisted Group Rings
In particular, we see that $\operatorname{\mathbf{Rep}}_{R}^{\hskip 0.81949pt\sigma}(G)$ is canonically identified with the full subcategory of $\operatorname{\mathbf{Mod}}_{R\rtimes_{\sigma}\hskip-0.81949ptG}$ where the underling $R$-module is free of finite rank.
@@ 3.1. Galois Extensions of a Field
$L\otimes_{K}L\rightarrow\prod_{\gamma\in\Gamma}L,\qquad a\otimes b\mapsto(ab(\gamma))_{\gamma\in\Gamma}$
@@ 3.2. Assumptions and Notation
Suppose momentarily that $L$ is a field and $G$ is a group that acts on $L$ through a group homomorphism $G\rightarrow\operatorname{Aut}(L)$ (as in the introduction). Setting $K\coloneqq L^{G}$, then the assumption that $L/K$ is a finite extension automatically implies that $L/K$ is Galois and that the induced map $G\rightarrow\operatorname{Gal}(L/K)$ is surjective (and therefore satisfies the assumptions of this section). Indeed, $K$ is the fixed field of a finite subgroup of $\operatorname{Aut}(L)$, so $L/K$ is Galois, and if $S\subset\operatorname{Gal}(L/K)$ denotes the image of $G$ then $L^{S}=L^{G}=K$, hence $S=\operatorname{Gal}(L/K)$ by the Galois correspondence.

### v3

@@ Abstract.
Let $G$ be a group acting on a field $L$, and suppose that $L/L^{G}$ is a finite extension. We show that the category of semilinear representations of $G$ over $L$ can be described in terms of the category of linear representations of $H$, the kernel of the map $G\rightarrow\operatorname{Aut}(L)$. When $G$ is finite and $L$ has characteristic 0 this provides a character theory for semilinear representations of $G$ over $L$, which recovers ordinary character theory when the action of $G$ on $L$ is trivial.
@@ 1. Introduction
$\rho(g)(\lambda\cdot v)=\sigma_{g}(\lambda)\cdot\rho(g)(v)\ \mbox{ for all }\ g\in G,\lambda\in L,v\in V.$
@@ 1.2. The Extension Problem
Similarly, if $W$ is the non-trivial one-dimensional representation of $C_{2}$ over ${\mathbb{Q}}(\sqrt{d})$, then to give a semilinear representation of $C_{4}$ over ${\mathbb{Q}}(\sqrt{d})$ extending $W$ is the same as giving an $a\in{\mathbb{Q}}(\sqrt{d})$ with
@@ 1.5. Group Action
1.6. Indecomposable Representations
The category $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ satisfies the Krull-Remak-Schmidt Theorem (Lemma 2.14), and therefore one is reduced to understanding $\operatorname{Ind}_{L}^{\rtimes}(G)$, the set of isomorphism classes of indecomposable semilinear representations of $G$ over $L$. We similarly write $\operatorname{Ind}_{L}(H)$ for the set of isomorphism classes of indecomposable representations of $H$ over $L$.
Our next result describes $\operatorname{Ind}_{L}^{\rtimes}(G)$ in terms of $\operatorname{Ind}_{L}(H)$ and its action of $\operatorname{Gal}(L/K)$. To ease notation, we set $\Gamma\coloneqq\operatorname{Gal}(L/K)$, and for $W\in\operatorname{Ind}_{L}(H)$ write $\Gamma_{W}$ for the stabiliser of $W$ in $\Gamma$.
Along with the restriction functor
@@ 1.6. Indecomposable Representations
Suppose that $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$ and $W\in\operatorname{Ind}_{L}(H)$ is a direct summand of $V|_{H}$. Then:
@@
    for some integer $m(V)\geq 1$ with $m(V)\mid|\Gamma_{W}|$.
  • $V$ is irreducible if and only if $W$ is irreducible.
  • $\operatorname{Ind}_{H}^{G}(W)\cong\operatorname{Ind}_{H}^{G}(\gamma*W)$ for any $\gamma\in\Gamma$.
  • Any $W\in\operatorname{Ind}_{L}(H)$ is a direct summand of $V|_{H}$ for a unique $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$.
  • This defines bijective correspondences
    $\operatorname{Ind}_{L}^{\rtimes}(G)\leftrightarrow\operatorname{Ind}_{L}(H)/\hskip 1.00006pt\Gamma,\qquad\operatorname{Irr}_{L}^{\rtimes}(G)\leftrightarrow\operatorname{Irr}_{L}(H)/\hskip 1.00006pt\Gamma.$
In particular, this gives a complete description of $\operatorname{Ind}_{L}^{\rtimes}(G)$, up to a knowledge of the numbers $m(V)$: for any orbit of $\operatorname{Gal}(L/K)$ on $\operatorname{Ind}_{L}(H)$, $m(V)$ copies of the orbit sum extends uniquely to an indecomposable semilinear representation of $G$ over $L$, and all elements of $\operatorname{Ind}_{L}^{\rtimes}(G)$ arise this way.
@@ 1.7. Semilinear Schur Index
For $W\in\operatorname{Ind}_{L}(H)$, set $m_{K}^{L}(W)\coloneqq m(V)$ for the unique corresponding $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$.
We refer to $m_{K}^{L}(W)$ as the semilinear Schur index of $W$, as these generalise the classical Schur indices in an appropriate sense (see Section 10). In light of Theorem B, these measure the failure for a $\Gamma$-fixed $W\in\operatorname{Ind}_{L}(H)$ to extend to a semilinear representation.
@@
$W\in\operatorname{Ind}_{L}(H)$ extends to a semilinear representation of $G$ over $L$ if and only if $\Gamma$ fixes $W$ and $m_{K}^{L}(W)=1$.
In Section 9 we establish the basic properties of the $m_{K}^{L}(W)$, including the fact that all $m_{K}^{L}(W)=1$ whenever $W\in\operatorname{Irr}_{L}(H)$ and $L$ is a finite field. In general, like their classical counterparts, the numbers $m_{K}^{L}(W)$ are hard to determine and often contain arithmetic information. For example, for $W$ as in Example 1.1, $m_{K}^{L}(W)\in\{1,2\}$, and $m_{K}^{L}(W)=1$ if and only if $x^{2}-dy^{2}=-1$ has a solution over $\mathbb{Q}$}.
@@ 1.8. Character Theory
When $G$ is finite and $L$ has characteristic 0, character theory provides a powerful and computationally efficient tool for completely describing the category $\operatorname{\mathbf{Rep}}_{L}(G)$.
@@
Suppose that $G$ is finite and $L$ has characteristic 0. If $V,W\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, then:
@@
More generally, when $G$ is finite, $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ is semisimple if and only if $|H|\in L^{\times}$ (Corollary 8.21), and in this context one can still use character theory to understand $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ (Section 12).
For example, this gives a complete description of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ using only characters when $L$ is a finite field, as in this case all $m_{K}^{L}(W)=1$.
As an example of how this character theory works, we give a complete classification of the semilinear representations of $S_{3}$ acting on a general degree two Galois extension $L/K$ via the natural projection $S_{3}\twoheadrightarrow\operatorname{Gal}(L/K)$. This is first done by hand in Section 5, in the case of characteristic 0, with explicit semilinear matrix computations. We then return to this example later in Section 13 to recover this classification more conceptually using characters and Theorems A and B.
@@ 1.10. Determination of Semilinear Schur Indices
By Theorem B, to understand $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ one is reduced to understanding the semilinear Schur indices of $W\in\operatorname{Ind}_{L}(H)$. This is hard in general.
@@ 1.10.1. Other Associated Linear Categories
One natural way one might try to glean information about the semilinear Schur indices $m_{K}^{L}(W)$ is to use other linear representation categories associated with $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$. Theorem B and the theory developed above can be viewed as a partial description of $\operatorname{Ind}_{L}^{\rtimes}(G)$ in terms of $\operatorname{Ind}_{L}(H)$, but there is another linear representation category that one might naturally consider, namely $\operatorname{\mathbf{Rep}}_{K}(G)$.
@@ 1.11. Structure of the Paper
In Section 4 we study those semilinear representations arising from matrix representations defined over $K$, and give an extended example of how these can be used in some cases to describe all semilinear representations of $G$ over $L$ in Section 5, which also serves to demonstrate how one works with semilinear representations explicitly.
@@
In Section 7 we show that the semilinear representation category $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, for $L$ a potentially disconnected finite Galois extension of $K$, is equivalent to $\operatorname{\mathbf{Rep}}_{F}^{\rtimes}(G_{e})$, where $F$ is the field defined by a connected component $\operatorname{Spec}(L)$, and $G_{e}$ is the stabiliser. This reduces general semilinear representation categories to the case when $L$ is a field, and is used in Section 8 to prove Theorems A and B.
In Section 8, we prove Theorems A and B. The key idea is to interpret the functors $(-)|_{H}$ and $\operatorname{Ind}_{H}^{G}$ between $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ and $\operatorname{\mathbf{Rep}}_{L}(H)$ as certain honest base change functors for the field extension $L/K$. This allows us to port many properties of base change to our setting, such as the preservation of semisimplicity, and how the functors $(-)|_{H}$ and $\operatorname{Ind}_{H}^{G}$ compose. Theorem A follows from Corollary 8.6, Remark 2.13, Remark 3.5, and Remark 8.9, and Theorem B follows from Remark 3.5 and Theorem 8.20. We also relate the integers $m(V)$ to the $K$-algebras $\operatorname{End}_{L\rtimes G}(V)$ in Proposition 8.29.
In Section 9 we define the semilinear Schur index of an object of $\operatorname{Ind}_{L}(H)$, and establish its basic properties, and in Section 10 we explain how the semilinear Schur indices specialise to the (classical) Schur indices in the case when $H\hookrightarrow G$ is split. We also highlight which properties the classical Schur indices satisfy which are not satisfied by all semilinear Schur indices.
In Section 11 we study the splitting behaviour of the $K$-division algebras $\operatorname{End}_{L\rtimes G}(V)$ for $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$, and consider base change for semilinear representations. When $K$ is a number field we use the local-global principle for $\operatorname{Br}(K)$ to prove a local-global principle for semilinear representations.
@@ 2. Semilinear Representations
$\rho(g)(\lambda\cdot v)=\sigma_{g}(\lambda)\cdot\rho(g)(v)\ \mbox{ for all }\ g\in G,\lambda\in R,v\in V.$
@@
Definition 2.5.
$\rho(g)(w)\in W\ \mbox{ for all }\ g\in G,w\in W.$
We say that $V$ is irreducible if $V$ is non-zero and 0 and $V$ are the only subrepresentations of $V$, and write $\operatorname{Irr}_{R}^{\hskip 0.81949pt\sigma}(G)$ for the set of isomorphism classes of irreducible objects of $\operatorname{\mathbf{Rep}}_{R}^{\hskip 0.81949pt\sigma}(G)$.
We say that $V$ is indecomposable if $V$ is non-zero and $V$ cannot be written as the direct sum of two non-zero subrepresentations, and write $\operatorname{Ind}_{R}^{\hskip 0.81949pt\sigma}(G)$ for the set of isomorphism classes of indecomposable objects of $\operatorname{\mathbf{Rep}}_{R}^{\hskip 0.81949pt\sigma}(G)$.
@@ 2.1.2. Semilinear Matrix Representations
Let $n\coloneqq\text{rank}_{R}(V)$, and fix a basis of the $R$-module $V$. For any $g\in G$, write $A_{g}\in M_{n\times n}(R)$ for the matrix of the linear map $\rho(g)\colon{}^{\sigma_{g}}V\rightarrow V$. The compatibility relations above in matrix form become
@@
We can give matrix descriptions of the constructions of Examples 2.3 and 2.4. The trivial representation has matrix representation by $(I_{1})_{g\in G}$, for $I_{1}$ the identity of $M_{1}(R)$, and if $V$ and $W$ in $\operatorname{\mathbf{Rep}}_{R}^{\hskip 0.81949pt\sigma}(G)$ have matrix representations $(A_{g})_{g\in G}$ and $(B_{g})_{g\in G}$ respectively, then
@@ 2.2. Semilinear Representations and Twisted Group Rings
In particular, we see that $\operatorname{\mathbf{Rep}}_{R}^{\hskip 0.81949pt\sigma}(G)$ is canonically identified with the full subcategory of $\operatorname{\mathbf{Mod}}_{R\rtimes_{\sigma}\hskip-0.81949ptG}$ where the underlying $R$-module is free of finite rank.
@@ 3.1. Galois Extensions of a Field
$L\otimes_{K}L\rightarrow\prod_{\gamma\in\Gamma}L,\qquad a\otimes b\mapsto(a\gamma(b))_{\gamma\in\Gamma}$
@@ 3.2. Assumptions and Notation
Suppose momentarily that $L$ is a field and $G$ is a group that acts on $L$ through a group homomorphism $G\rightarrow\operatorname{Aut}(L)$ (as in the introduction). Setting $K\coloneqq L^{G}$, then the assumption that $L/K$ is a finite extension automatically implies that $L/K$ is Galois and that the induced map $G\rightarrow\operatorname{Gal}(L/K)$ is surjective (and therefore satisfies the assumptions of this section). Indeed, $K$ is the fixed field of a subgroup of the finite group $\operatorname{Aut}(L/K)$, so $L/K$ is Galois, and if $S\subset\operatorname{Gal}(L/K)$ denotes the image of $G$ then $L^{S}=L^{G}=K$, hence $S=\operatorname{Gal}(L/K)$ by the Galois correspondence.